OpenStudy (omgthatgirl):

What is the value of q? A: 58 B: 122 C: 180 D: 116

OpenStudy (whpalmer4):

Probably one of those, hard to choose without more details :-)

OpenStudy (whpalmer4):

Okay, how much will the four angles sum to?

OpenStudy (omgthatgirl):

360

OpenStudy (whpalmer4):

do we know anything else about the angles of a quadrilateral inscribed in a circle?

OpenStudy (omgthatgirl):

No

OpenStudy (whpalmer4):

well, it turns out that not only do they sum to 360 degrees, but opposite angles are supplementary, so they sum to 180 degrees.

OpenStudy (whpalmer4):

From that, can you figure out the values of p and q?

OpenStudy (omgthatgirl):

i think so

OpenStudy (whpalmer4):

operators are standing by to check your answer :-)

OpenStudy (omgthatgirl):

is it d?

OpenStudy (whpalmer4):

p is complementary with a 90 degree angle, so p + 90 = 180, so p = 180-90 = 90 we know that q and 58 degrees are complementary, so q + 58 = 180 or 180 -58 = q which means q =

OpenStudy (omgthatgirl):

122

OpenStudy (omgthatgirl):

thank you

OpenStudy (whpalmer4):

take a look at this web page: http://www.mathopenref.com/quadrilateralinscribedangles.html it has a cool gadget for demonstrating this property

OpenStudy (whpalmer4):

you can even drag it into position to check your answer :-)